# Dual Modalities of Text: Visual and Textual Generative Pre-training

### v2

@@ 4.3 Analysis
Font Transfer Analysis We extend to examining the adaptability of PixelGPT to diverse font styles during fine-tuning. We employed three distinct fonts for rendering the data: GoNotoCurrent, which was utilized during pre-training; NotoSerif-Regular, a font stylistically akin to GoNotoCurrent; and JournalDingbats1, a font that renders text as distinct image-based symbols, markedly divergent from the others. The adaptability was tested across five datasets from the GLUE benchmark—CoLA, STS-B, MRPC, RTE, and WNLI. As depicted in Figure 6, the performance of PixelGPT remained stable across different fonts for all selected datasets barring CoLA. Notably, even when fine-tuned with data rendered in JournalDingbats1, which bears little resemblance to the pre-training font, the results demonstrated a commendable degree of resilience, indicating that the pixel pre-training is robust to generalize across significantly varied visual representations.
@@
Impact Analysis of Color Retention Unlike PIXAR that renders text as binary images, PixelGPT employs RGB-rendered data, retaining richer informational content. We evaluated the performance of these rendering approaches on HatemojiBuild dataset Kirk et al. (2022), designed for detecting online hate speech conveyed through emojis. Table 5 presents our findings, where the RGB-rendered data fine-tuning significantly outperforms its grayscale counterpart. This performance enhancement can be attributed to the model’s capacity to utilize color cues within emojis, which are critical for inferring the emotional context of sentences. For a more detailed illustration, Figure 7 provides specific examples where color retention has improved model interpretability.
@@ 5 Conclusion and Future Work
Environmental Impact
The training of large-scale models is resource-intensive and has a significant environmental footprint. We must consider sustainable practices in model training, including optimizing computational efficiency and exploring energy-efficient hardware to reduce the overall carbon emissions associated with our research.
@@
Continual Monitoring and Evaluation
Post-deployment monitoring and ongoing evaluation of the model’s performance and societal impact are crucial. This process helps ensure the model adapts to changes over time and continues to operate within the ethical boundaries set forth by evolving standards and expectations.
By addressing these ethical considerations, we aim to promote responsible research and application of advanced machine learning techniques in language processing, contributing positively to the field and society at large.
@@ E.2 Fine-tuning Setting
We searched fine-tuning hyperparameters for each dataset in GLUE and two XNLI settings for PixelGPT, MonoGPT, DualGPT and TextGPT, respectively. Table 11 shows the searched hyperparameters and values. We present the best searched results for GLUE in Table 14 and Table 14 and for translate-train-all and cross-lingual transfer settings on XNLI in Table 14. During the hyperparameter searching, we found that using a larger batch size to fine-tune the generative pixel-based pre-training model improves training stability and achieves better results on some datasets. For a detailed analysis, see § 4.
@@ F.2 Image-based Baselines
PIXAR
PIXAR Tai et al. (2024) is a pixel-based pre-trained autoregressive language model on binary text images. PIXAR operates directly on pixel representations of text, enabling it to perform free-form generative tasks without the constraints of a predefined vocabulary. The model architecture employs a decoder-only setup, similar to GPT-like models but adapted to handle pixel input. PIXAR represents a significant advancement in the field by enabling text generation in images and leveraging a novel adversarial pre-training stage to enhance the clarity and accuracy of generated text.

### v3

@@ A Large Batch Size Improves Stable Training
Font Transfer Analysis
We extend to examining the adaptability of PixelGPT to diverse font styles during fine-tuning. We employed three distinct fonts for rendering the data: GoNotoCurrent, which was utilized during pre-training; NotoSerif-Regular, a font stylistically akin to GoNotoCurrent; and JournalDingbats1, a font that renders text as distinct image-based symbols, markedly divergent from the others. The adaptability was tested across five datasets from the GLUE benchmark—CoLA, STS-B, MRPC, RTE, and WNLI. As depicted in Figure 6, the performance of PixelGPT remained stable across different fonts for all selected datasets barring CoLA. Notably, even when fine-tuned with data rendered in JournalDingbats1, which bears little resemblance to the pre-training font, the results demonstrated a commendable degree of resilience, indicating that the pixel pre-training is robust to generalize across significantly varied visual representations.
@@
Impact Analysis of Color Retention
Unlike previous that renders text as grayscale or binary images, PixelGPT employs RGB-rendered data, retaining richer informational content. We evaluated the performance of these rendering approaches on HatemojiBuild dataset Kirk et al. (2022), designed for detecting online hate speech conveyed through emojis. Table 6 presents our findings, where the RGB-rendered data fine-tuning significantly outperforms its grayscale counterpart. This performance enhancement can be attributed to the model’s capacity to utilize color cues within emojis, which are critical for inferring the emotional context of sentences. For a more detailed illustration, Figure 7 provides specific examples where color retention has improved model interpretability.
@@ 5 Conclusion and Future Work
Acknowledgements
We would like to thank all anonymous reviewers for their insightful and constructive feedback.
@@ Acknowledgements
Extended Evaluation on Text Generation
One limitation of our approach is related to generation tasks. Since the model’s input and output are image patches, directly obtaining text outputs requires an additional OCR postprocessing step. This introduces an additional layer of complexity and potential error. We plan to address this in future work, exploring more integrated solutions for text generation tasks.
@@ E.2 Fine-tuning Setting
We searched fine-tuning hyperparameters for each dataset in GLUE and two XNLI settings for PixelGPT, MonoGPT, DualGPT and TextGPT, respectively. Table 12 shows the searched hyperparameters and values. We present the best searched results for GLUE in Table 15 and Table 15 and for translate-train-all and cross-lingual transfer settings on XNLI in Table 15. During the hyperparameter searching, we found that using a larger batch size to fine-tune the generative pixel-based pre-training model improves training stability and achieves better results on some datasets. For a detailed analysis, refer to § 4.3.
@@ G.4 Comparison on XNLI under Translate-Train-All Settings
G.5 Benefits of Pixel-based Models
Our pixel-based method offers significant advantages:
  1. 1.
    Tokenization-Free: Pure pixel-based training (w/o texts) eliminates the need for tokenization, thereby removing the vocabulary bottleneck problem, which is critical for handling diverse linguistic constructs and scaling effectively to multilingual contexts.
  2. 2.
    Rich Visual Representation: Leverages the rich information content of real-valued RGB images, capturing nuances that text-based tokenization may miss.
  3. 3.
    Modality Interplay: Demonstrates the potential for effective integration of visual and textual data, enhancing the overall model performance in language understanding tasks.
While all language models with pixel-based modalities currently match or slightly underperform compared to text modality models, the potential for scaling and the removal of tokenization challenges present a compelling case for further development and research in this area.